OpenStudy (anonymous):

AVOGADRO LAW at a certain T and P, a balloon w/ 10.0g of oxygen has a volume of 7.00L. what is the volume after 5.00g of oxygen is added to the balloon?

OpenStudy (anonymous):

Using ratios 10/ 7 = 1.4285 so 15 / x = 1.4285 x = 15 / 1.4285 x = 10.5